# CODA: Repurposing Continuous VAEs for Discrete Tokenization

### v1

@@ Abstract
Discrete visual tokenizers transform images into a sequence of tokens, enabling token-based visual generation akin to language models. However, this process is inherently challenging, as it requires both compressing visual signals into a compact representation and discretizing them into a fixed set of codes. Traditional discrete tokenizers typically learn the two tasks jointly, often leading to unstable training, low codebook utilization, and limited reconstruction quality. In this paper, we introduce CODA (COntinuous-to-Discrete Adaptation), a framework that decouples compression and discretization. Instead of training discrete tokenizers from scratch, CODA adapts off-the-shelf continuous VAEs—already optimized for perceptual compression—into discrete tokenizers via a carefully designed discretization process. By primarily focusing on discretization, CODA ensures stable and efficient training while retaining the strong visual fidelity of continuous VAEs. Empirically, with $\mathbf{6\times}$ less training budget than standard VQGAN, our approach achieves a remarkable codebook utilization of 100% and notable reconstruction FID (rFID) of $\mathbf{0.43}$ and $\mathbf{1.34}$ for 8×8$\times$8 × and 16×16$\times$16 × compression on ImageNet 256$\times$ 256 benchmark.
@@ 1 Introduction
The field of AI-generated content (AIGC) has witnessed significant progress in recent years. In natural language processing, text generation has been mainly unified by the paradigm of next discrete token prediction [44, 34, 6, 1]. This contrasts with developments in computer vision, where the debate between continuous and discrete generation paradigms has yet to reach a conclusion. While continuous diffusion models dominate the field [36, 31, 30, 5], discrete token-based approaches [7, 8, 29, 40, 41] are increasingly gaining interest due to their computational efficiency [41, 28] and promising potential for unifying tasks across language modeling and multimodal understanding [52, 46, 51, 49].
@@
Building upon this conceptual framework, we craft a set of CODA tokenizers, instantiated through systematic adaptations on the 16×16$\times$16 × compression MAR [21] and 8×8$\times$8 × compression FLUX [5] VAEs. Quantitative evaluations demonstrate that CODA achieves competitive performance of $1.34$ and $0.43$ rFID on the ImageNet dataset, respectively, while maintaining full codebook utilization. Moreover, our proposed disentangling leads to high computational efficiency in training, enabling a reduction of 6×6$\times$6 × in training compute compared to representative prior VQGAN training settings. Empirical results also demonstrate that, when integrated with the CODA tokenizer, representative discrete token-based generation methods, i.e. MaskGIT [7] can be significantly improved to achieve overall competitive performance against main stream continuous and discrete generation paradigms.
@@ 2.1 Continuous and Discrete Image Synthesis
Continuous approaches to image synthesis, i.e. latent diffusion models [36, 31, 30, 5] operate on a latent space sampled from a Variational Autoencoder (VAE) [18]. Initially, images are transformed from pixel space into a compressed, continuous latent space using a VAE, where they undergo iterative denoising via diffusion models. This process involves progressively refining the noisy latent representations through multiple denoising steps, leveraging trained neural networks to construct high-level semantic features and fine details. Once denoising is complete, the latents are decoded back into pixel space to generate the final image. This approach has been proven highly successful, enabling the generation of high-quality, photorealistic images.
Discrete approaches to image synthesis utilize discrete token representations as means for generative modeling. Images in pixel space are first transformed using image tokenizers  [43, 11] into a discrete token space. Multiple paradigms build upon this discrete token space and iteratively generate the predicted token combination, after which the predicted tokens are decoded back to pixel space. Autoregressive  [19, 40, 55] models draw inspiration from language modeling, treating images as a sequence of discrete tokens and generating them by iteratively predicting the next token in the sequence. MaskGITs  [7, 8, 29] models leverage the bi-directional nature of images by enabling parallel decoding. These models simultaneously unmask a combination of discrete tokens at each forward step, improving the efficiency and speed of the generation process. By avoiding the sequential dependencies inherent in autoregressive models, MaskGIT approaches can generate images more rapidly while maintaining competitive performance. Visual autoregressive  [41] models are inspired by the coarse-to-fine approach in image synthesis and adopt next-scale prediction to predict a combination of discrete tokens from the same scale at each step. Discrete token-based generation is gradually gaining popularity and becoming a promising approach to visual synthesis, as it is more efficient in terms of inference speed, aligns and adapts to similar paradigms in natural language processing  [32, 33, 42], multi-modal understanding [22, 46, 52] and Embodied AI [50].
@@ 4 CODA Tokenizer
Continuous VAE [21] - $0.69$ $24.9$ $71.6$
Vector Quant. $0.181$ $41.52$ $18.9$ $46.1$ $49\%$
+ Residual Quant. $0.094$ $14.23$ $20.6$ $53.3$ $51\%$
+ Attention-based Quant. $0.065$ $7.06$ $21.4$ $57.1$ $100\%$
+ Adapt VAE $0.001$ $1.34$ $22.2$ $60.2$ $100\%$
Table 1: Summary of our main ablation results in building the CODA tokenizer. We start from an off-the-shelf continous VAE [21] (marked in gray) and progressively introduce design components for effective tokenization. The codebook size is fixed to 65536. Quant: quantization. Util: utilization of codebook.
@@
To analyze the reason behind the performance drop, we first visualize the continuous latent space of the pretrained VAE and the discretized space after applying vector quantization in Figure 3 (a) and (b). Results indicate a substantial information loss during the VQ approximation process. More specifically, continuous features span across large domains in the latent space, covering not only high-density areas in the center but also marginal regions that are less populated. In contrast, discrete features quantized by VQ occupy only sparse points in the latent space. This outcome is not surprising, though, as the number of features that VQ can represent is inherently limited by the number of code embeddings in codebook. Though we have alreadly employed an extensive codebook size of $65536$, exceeding that of many representative VQ tokenizers [11, 53, 40, 48], the representational capacity still remains far from sufficient to capture the full diversity of the densely populated continuous latent space.
@@
This design introduces a learnable mechanism for assigning features to discrete codes with enhanced sparsity and improved training dynamics. Attention-based quantization frames discretization as a retrieval task, where the quantizer learns to retrieve the most appropriate index from a large codebook. Under this framework, different codes compete and suppress each other when calculating softmax attention scores, enforcing sparse assignment. This is confirmed by the visualization of confidence in Figure 5(b) (b). Additionally, this approach delivers a push effect along with pull during training: the chosen code index is pulled towards the clustering centroids, reducing quantization error, while irrelevant codes are pushed away to other regions of the latent space. This new dynamic ensures more effective codebook usage and comprehensive coverage of the latent space, as visualuization of training dynamics in Figure 6 (b) illustrate. Quantitative results in Tab. 1 shows that our the attention-based assignment achieves a remarkable 100% codebook utilization, and significantly improves the reconstruction FID from $14.23$ to $7.06$.
[…]
VQGAN [11] $16$ $1024$ $7.94$ $19.4$ $50.0$ $44$ $300$K
VQGAN [11] $16$ $16384$ $4.98$ $19.9$ $51.0$ $5.9$ $300$K
DF-VQGAN [27] $16$ $12288$ $5.16$
DQVAE [15] $16$ $1024$ $4.08$ $250$K
DiVAE [39] $16$ $16384$ $4.07$
VQGAN-LC  [59] $16$ $16384$ $3.01$ $23.2$ $56.4$ $99$ $100$K
VQGAN-LC  [59] $16$ $100000$ $2.62$ $23.8$ $58.9$ $99$ $100$K
LlamaGen  [40] $16$ $16384$ $2.19$ $20.8$ $67.5$ $97$ $200$K
RQVAE [19] - $16384$ $2.69$ $250$K
BAE [47] $16$ $65536$ $3.32$ $2000$K
MaskBit [48] $16$ $16384$ $1.61$ $1350$K
CODA MAR $16$ $16384$ $1.43$ $22.0$ $59.4$ $100$ 50K
CODA MAR $16$ $65536$ 1.34 $22.2$ $60.2$ $100$ 50K
[…]
VQGAN [11] $8$ $16384$ $1.14$ $23.4$ $67.0$ $5.4$
ViT-VQGAN [53] $8$ $8192$ $1.28$
DF-VQGAN [27] $8$ $8192$ $1.38$
DiVAE [39] $8$ $16384$ $1.28$
OmniTokenizer [45] $8$ $8192$ $1.11$ $24.0$ $75.2$ $100$
CODA FLUX $8$ $65536$ 0.43 25.9 77.1 $100$ 50K
[…]
@@ 5.1 Experiment Setup
We conduct our experiments on the widely used large-scale ImageNet dataset [9], specifically utilizing images at a resolution of $256\times 256$. Our study involves both training our tokenizer and evaluating generative models using this dataset. To assess the reconstruction quality of our approach, we employ multiple evaluation metrics, including reconstruction FID (rFID), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index Measure (SSIM). These metrics provide a comprehensive evaluation of both perceptual and pixel-level reconstruction accuracy.
@@
Leveraging the continuous latent space of pretrained VAEs, we utilize and freeze the encoder-decoder weights from two standard continuous VAE: a 16×16$\times$16 × MAR [21] VAE and a 8×8$\times$8 × FLUX VAE [5]. To adapt the continuous encoder-decoder to discrete codes, we implement LoRA [14] modules exclusively on the convolution blocks and finetune the added modules along with the quantizer during training. In practice, we adopt a LoRA rank of $32$ for both encoder and decoder. We follow a main stream implementation of residual quantization introduced in  [41] and set the quantization level to $10$. All models are trained with an accumulated batch size of $512$ for $10$ epochs on the ImageNet $256\times 256$ dataset with a $45e-$ 4 learning rate.
@@
MaskGIT models are trained for $500$K steps with an accumulated batch size of $2048$ on the ImageNet $256\times 256$ dataset with $44e-$ 4 learning rate. In addition to training on the standard tokenized representation, we introduce learnable positional embeddings specific to each quantization level. This modification enhances the model’s ability to differentiate between tokens from different levels of the residual quantization hierarchy, ultimately improving the decoding process.
@@ 5.2 Main Results
We present the reconstruction results in Tab. 2. There are three key observations from our findings: First, our CODA tokenizer achieves competitive performance with $0.43$ and $1.34$ rFID in 8×8$\times$8 × and 16×16$\times$16 × reduction settings, respectively, significantly outperforming the corresponding baseline models in terms of reconstruction quality and fidelity. Second, our tokenizer achieves a high utilization rate of $100\%$ with a large codebook of $65536$, compared to VQGAN’s utilization rate of $5.9\%$ with $16384$ codes. This demonstrates that the combined approach of training in a pretrained continuous latent space and employing attention-based sparse quantization mechanisms prevents codebook collapse and encourages high utilization, providing the potential for scaling to even larger scopes. Finally, our approach requires substantially less training computation than all baseline methods, achieving a 6×6$\times$6 × speedup compared to the representative VQGAN training recipe. Additional qualitative visualizations and comparisons are provided in Figure 7 (a).
@@
ADM [10] Diff. $554$M $250$ $10.94$
LDM-4-G [36] Diff. $400$M $250$ $3.60$
DiT-L/2 [30] Diff. $458$M $250$ $5.02$
UViT-L/2 [2] Diff. $287$M $50$ $3.40$
LDM-4-G [36] Diff. $400$M $8$ $4.56$
DiT-XL/2 [30] Diff. $675$M $8$ $5.18$
RQVAE-GPT [19] AR $480$M $256$ $15.7$
VQGAN-LC-GPT [59] AR $404$M $256$ $15.4$
ViT-VQGAN-GPT [53] AR $650$M $1024$ $8.81$
RQ-Transformer [19] AR $3.8$B $64$ $7.55$
LlamaGen-L-384 [40] AR $343$M $576$ $3.07$
VAR [41] VAR $310$M $10$ $3.30$
MaskGIT [7] Mask. $227$M $12$ $4.92$
MaskGIT-FSQ [26] Mask. $225$M $12$ $4.53$
MAGE [20] Mask. $230$M $20$ $6.93$
ENAT [28] Mask. $219$M $8$ $3.53$
CODA MaskGIT-L Mask. $195$M $8$ $3.17$
CODA MaskGIT-L Mask. $195$M $32$ 2.66
@@
In Tab. 3, We compare our approach with leading image generation paradigms, including continuous diffusion-based methods, discrete MaskGIT, autoregressive models, and visual autoregressive models. Our key observations are as follows: 1) Compared to other MaskGIT-based models, our approach achieves a lower FID of $3.17$, compared to $4.92$ for MaskGIT and $4.53$ for MaskGIT-FSQ, demonstrating enhanced generation quality delivered by a more effective tokenization strategy. 2) Our model demonstrates strong overall performance relative to other main stream generation paradigms, with an FID of $2.66$ compared to $3.60$ for LDM and $3.07$ for LlamaGen. 3) Our approach achieves high-quality generation results comparable to continuous methods while maintaining the superior efficiency of discrete models, requiring fewer sampling steps. Additional qualitative results of generated images from our approach are demonstrated in Figure 7 (b).

### v2

@@ Abstract
Discrete visual tokenizers transform images into a sequence of tokens, enabling token-based visual generation akin to language models. However, this process is inherently challenging, as it requires both compressing visual signals into a compact representation and discretizing them into a fixed set of codes. Traditional discrete tokenizers typically learn the two tasks jointly, often leading to unstable training, low codebook utilization, and limited reconstruction quality. In this paper, we introduce CODA (COntinuous-to-Discrete Adaptation), a framework that decouples compression and discretization. Instead of training discrete tokenizers from scratch, CODA adapts off-the-shelf continuous VAEs—already optimized for perceptual compression—into discrete tokenizers via a carefully designed discretization process. By primarily focusing on discretization, CODA ensures stable and efficient training while retaining the strong visual fidelity of continuous VAEs. Empirically, with $\mathbf{6\times}$ less training budget than standard VQGAN, our approach achieves a remarkable codebook utilization of 100% and notable reconstruction FID (rFID) of $\mathbf{0.43}$ and $\mathbf{1.34}$ for $8\times$ and $16\times$ compression on ImageNet $\times 256$ benchmark.
@@ 1 Introduction
The field of AI-generated content (AIGC) has witnessed significant progress in recent years. In natural language processing, text generation has been mainly unified by the paradigm of next discrete token prediction [45, 35, 6, 1]. This contrasts with developments in computer vision, where the debate between continuous and discrete generation paradigms has yet to reach a conclusion. While continuous diffusion models dominate the field [37, 32, 31, 5], discrete token-based approaches [7, 8, 29, 41, 42] are increasingly gaining interest due to their computational efficiency [42, 28, 30] and promising potential for unifying tasks across language modeling and multimodal understanding [53, 47, 52, 50].
@@
Building upon this conceptual framework, we craft a set of CODA tokenizers, instantiated through systematic adaptations on the $16\times$ compression MAR [21] and $8\times$ compression FLUX [5] VAEs. Quantitative evaluations demonstrate that CODA achieves competitive performance of $1.34$ and $0.43$ rFID on the ImageNet dataset, respectively, while maintaining full codebook utilization. Moreover, our proposed disentangling leads to high computational efficiency in training, enabling a reduction of $6\times$ in training compute compared to representative prior VQGAN training settings. Empirical results also demonstrate that, when integrated with the CODA tokenizer, representative discrete token-based generation methods, i.e. MaskGIT [7] can be significantly improved to achieve overall competitive performance against main stream continuous and discrete generation paradigms.
@@ 2.1 Continuous and Discrete Image Synthesis
Continuous approaches to image synthesis, i.e. latent diffusion models [37, 32, 31, 5] operate on a latent space sampled from a Variational Autoencoder (VAE) [18]. Images are first mapped from pixel space into this compressed representation and then subjected to iterative denoising via diffusion models. At each step, a trained neural network refines the latent by removing noise and rebuilding semantic structure and fine details. Once denoising is complete, the latents are decoded back into pixel space to generate the final image. This approach has been proven highly successful, enabling the generation of high-quality, photorealistic images.
Discrete approaches to image synthesis utilize discrete token representations as means for generative modeling. Images in pixel space are first transformed using image tokenizers  [44, 11] into discrete tokens. Multiple paradigms build upon this discrete space and iteratively generate the predicted token combination, after which the tokens are decoded back to pixel space. Autoregressive  [19, 41, 56] models draw inspiration from language modeling, treating images as a sequence of discrete tokens and generating them by iteratively predicting the next token in the sequence. MaskGITs  [7, 8, 29] models leverage the bi-directional nature of images by enabling parallel decoding. These models simultaneously unmask a combination of discrete tokens at each forward step, improving the efficiency and speed of the generation process. Visual autoregressive  [42] models are inspired by the coarse-to-fine approach in image synthesis and adopt next-scale prediction to predict a combination of discrete tokens from the same scale at each step. Discrete token-based generation is gradually gaining popularity and becoming a promising approach to visual synthesis, as it is more efficient in terms of inference speed, aligns and adapts to similar paradigms in natural language processing  [33, 34, 43], multi-modal understanding [22, 47, 53] and Embodied AI [51].
@@ 4 CODA Tokenizer
Continuous VAE [21] - $0.69$ $24.9$ $0.716$ -
Vector Quant. $0.181$ $41.52$ $18.9$ $0.461$ $49\%$
+ Residual Quant. $0.094$ $14.23$ $20.6$ $0.533$ $51\%$
+ Attention-based Quant. $0.065$ $7.06$ $21.4$ $0.571$ $100\%$
+ Adapt VAE $0.001$ $1.34$ $22.2$ $0.602$ $100\%$
Table 1: Summary of our main ablation results in building the CODA tokenizer. We start from an off-the-shelf continuous VAE [21] (marked in gray) and progressively introduce design components for effective tokenization. The codebook size is fixed to 65536. Quant: quantization. Util: utilization of codebook.
@@
To analyze the reason behind the performance drop, we first visualize the continuous latent space of the pretrained VAE and the discretized space after applying vector quantization in Figure 3 (a) and (b). Results indicate a substantial information loss during the VQ approximation process. More specifically, continuous features span across large domains in the latent space, covering not only high-density areas in the center but also marginal regions that are less populated. In contrast, discrete features quantized by VQ occupy only sparse points in the latent space. This outcome is not surprising, though, as the number of features that VQ can represent is inherently limited by the number of code embeddings in codebook. Though we have already employed an extensive codebook size of $65536$, exceeding that of many representative VQ tokenizers [11, 54, 41, 49], the representational capacity still remains far from sufficient to capture the full diversity of the densely populated continuous latent space.
@@
This design introduces a learnable mechanism for assigning features to discrete codes with enhanced sparsity and improved training dynamics. Attention-based quantization frames discretization as a retrieval task, where the quantizer learns to retrieve the most appropriate index from a large codebook. Under this framework, different codes compete and suppress each other when calculating softmax attention scores, enforcing sparse assignment. This is confirmed by the visualization of confidence in Figure 5(b) (b). Additionally, this approach delivers a push effect along with pull during training: the chosen code index is pulled towards the clustering centroids, reducing quantization error, while irrelevant codes are pushed away to other regions of the latent space. This new dynamic ensures more effective codebook usage and comprehensive coverage of the latent space, as visualization of training dynamics in Figure 6 (b) illustrate. Quantitative results in Tab. 1 shows that our the attention-based assignment achieves a remarkable 100% codebook utilization, and significantly improves the reconstruction FID from $14.23$ to $7.06$.
Method Ratio Codebook Size rFID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Utilization ($\%$) Training (iter)
VQGAN [11] $16$ $1024$ $7.94$ $19.4$ $0.500$ $44$ 300300K
VQGAN [11] $16$ $16384$ $4.98$ $19.9$ $0.510$ $5.9$ 300300K
DF-VQGAN [27] $16$ $12288$ $5.16$ - - - -
DQVAE [15] $16$ $1024$ $4.08$ - - - 250250K
DiVAE [40] $16$ $16384$ $4.07$ - - - -
VQGAN-LC  [60] $16$ $16384$ $3.01$ $23.2$ $0.564$ $99$ 100100K
VQGAN-LC  [60] $16$ $100000$ $2.62$ $23.8$ $0.589$ $99$ 100100K
LlamaGen  [41] $16$ $16384$ $2.19$ $20.8$ $0.675$ $97$ 200200K
RQVAE [19] - $16384$ $2.69$ - - - 250250K
BAE [48] $16$ $65536$ $3.32$ - - - 20002000K
MaskBit [49] $16$ $16384$ $1.61$ - - - 13501350K
CODA MAR $16$ $16384$ $1.43$ $22.0$ $0.594$ $100$ 50K
CODA MAR $16$ $65536$ 1.34 $22.2$ $0.602$ $100$ 50K
MAR [21] VAE $16$ - $0.69$ $24.9$ $0.716$ - -
VQGAN [11] 88 $16384$ $1.14$ $23.4$ $0.670$ $5.4$ -
ViT-VQGAN [54] 88 $8192$ $1.28$ - - - -
DF-VQGAN [27] 88 $8192$ $1.38$ - - - -
DiVAE [40] 88 $16384$ $1.28$ - - - -
OmniTokenizer [46] 88 $8192$ $1.11$ $24.0$ $0.752$ $100$ -
CODA FLUX 88 $65536$ 0.43 25.9 0.771 $100$ 50K
FLUX [5] VAE 88 - $0.17$ $31.1$ $0.903$ - -
@@ 5.1 Experiment Setup
We conduct our experiments on the ImageNet dataset [9], specifically utilizing images at a resolution of $256\times 256$. Our study involves both training our tokenizer and evaluating generative models using this dataset. To assess the reconstruction quality of our approach, we employ multiple evaluation metrics, including reconstruction FID (rFID), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index Measure (SSIM). These metrics provide a comprehensive evaluation of both perceptual and pixel-level reconstruction accuracy.
@@
Leveraging the continuous latent space of pretrained VAEs, we utilize and freeze the encoder-decoder weights from two standard continuous VAE: a $16\times$ MAR [21] VAE and a $8\times$ FLUX VAE [5]. To adapt the continuous encoder-decoder to discrete codes, we implement LoRA [14] modules exclusively on the convolution blocks and finetune the added modules along with the quantizer during training. In practice, we adopt a LoRA rank of $32$ for both encoder and decoder. We follow a main stream implementation of residual quantization introduced in  [42] and set the quantization level to $10$. All models are trained with an accumulated batch size of $512$ for $10$ epochs on the ImageNet $256\times 256$ dataset with a 5e45e-4 learning rate.
@@
MaskGIT models are trained for 500500K steps with an accumulated batch size of $2048$ on the ImageNet $256\times 256$ dataset with 4e44e-4 learning rate. In addition to training on the standard tokenized representation, we introduce learnable positional embeddings specific to each quantization level. This modification enhances the model’s ability to differentiate between tokens from different levels of the residual quantization hierarchy, ultimately improving the decoding process.
@@ 5.2 Main Results
We present the reconstruction results in Tab. 2. There are three key observations from our findings: First, our CODA tokenizer achieves competitive performance with $0.43$ and $1.34$ rFID in $8\times$ and $16\times$ reduction settings, respectively, significantly outperforming corresponding baselines in terms of reconstruction quality and fidelity. Second, our tokenizer achieves a high utilization rate of $100\%$ with a large codebook of $65536$, compared to VQGAN’s utilization rate of $5.9\%$ with $16384$ codes. This demonstrates that the combined approach of training in a pretrained continuous latent space and employing attention-based sparse quantization mechanisms prevents codebook collapse and encourages high utilization, providing the potential for scaling to even larger scopes. Finally, our approach requires substantially less training computation than all baseline methods, achieving a $6\times$ speedup compared to the representative VQGAN training recipe. Additional qualitative visualizations and comparisons are provided in Figure 7 (a).
@@
ADM [10] Diff. 554554M $250$ $10.94$
LDM-4-G [37] Diff. 400400M $250$ $3.60$
DiT-L/2 [31] Diff. 458458M $250$ $5.02$
UViT-L/2 [2] Diff. 287287M $50$ $3.40$
LDM-4-G [37] Diff. 400400M 88 $4.56$
DiT-XL/2 [31] Diff. 675675M 88 $5.18$
RQVAE-GPT [19] AR 480480M $256$ $15.7$
VQGAN-LC-GPT [60] AR 404404M $256$ $15.4$
ViT-VQGAN-GPT [54] AR 650650M $1024$ $8.81$
RQ-Transformer [19] AR 3.83.8B $64$ $7.55$
LlamaGen-L-384 [41] AR 343343M $576$ $3.07$
VAR [42] VAR 310310M $10$ $3.30$
MaskGIT [7] Mask. 227227M $12$ $4.92$
MaskGIT-FSQ [26] Mask. 225225M $12$ $4.53$
MAGE [20] Mask. 230230M 2020 $6.93$
ENAT [28] Mask. 219219M 88 $3.53$
CODA MaskGIT-L Mask. 195195M 88 $3.17$
CODA MaskGIT-L Mask. 195195M $32$ 2.66
@@
In Tab. 3, We compare our approach with leading image generation paradigms, including continuous diffusion-based methods, discrete MaskGIT, autoregressive models, and visual autoregressive models. Our key observations are as follows: 1) Compared to other MaskGIT-based models, our approach achieves a lower FID of $3.17$, compared to $4.92$ for MaskGIT and $4.53$ for MaskGIT-FSQ, demonstrating enhanced generation quality delivered by a more effective tokenization strategy. 2) Our model demonstrates strong overall performance relative to other main stream generation paradigms, with an FID of $2.66$ compared to $3.60$ for LDM and $3.07$ for LlamaGen. 3) Our approach achieves high-quality generation results comparable to continuous methods while maintaining the superior efficiency of discrete models, requiring fewer sampling steps. Additional quantitative text-to-image generation results and qualitative results of generated images from our approach are demonstrated in  Tab. 5 and  Figure 7 (b) respectively.
[…]
Table 5: Text-to-image generation on MS-COCO.